A train travels a distance of 48 miles. It took 34 minutes. What was the average speed of the train?

Found 2 solutions by Alan3354, TimothyLamb:
Answer by Alan3354(69443)

A train travels a distance of 48 miles. It took 34 minutes. What was the average speed of the train?
===============================
d = r*t
r = d/t

Answer by TimothyLamb(4379)

s = d / t
48 / 34 = 1.41 mi/min
Answer:
1.41 mi/min
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php